OpenStudy (anonymous):

Identify the solutions of 2x2 + 13x + 6 = 0. x = 1 over 2, x = 6 x = −1 over 2, x = −6 x = −3 over 2, x = −2 x = 3 over 2, x = 2

OpenStudy (bradely):

2x^2 + 13x + 6 = 0 2x^2 + 12x +x+ 6 = 0 2x(x+6)+1(x+6)=0 (x+6)(2x+1)=0 x=-6 or -1/2 Option 2 is correct Source: http://www.mathskey.com/question2answer/